# Thermal Hall transport in Kitaev spin liquids

### v1

@@ I Introduction
where $k_{\mathrm{B}}$ and $\hbar$ denote the Boltzmann and Dirac constants, respectively. Here, $\nu$ takes ±$\pm 1$ depending on the direction of the magnetic field. The coefficient of $1/2$ is the consequence of the fractionalization; each Majorana fermion carries half the degrees of a conventional Dirac fermion. Thus, this quantized value is half of that for electrons with the same band topology, and hence, it is called the half-integer quantization. Detecting this peculiar behavior offers direct evidence of the Majorana fermions in the Kitaev QSL.
@@
For the Kitaev model without other additional interactions, efficient numerical simulation methods have been developed for studying thermodynamic properties and spin dynamics at finite temperatures, utilizing the Majorana fermion representation [49, 50, 51, 52, 53]. This approach explicitly leverages the emergent quasiparticle picture in fractionalization. Although these methods are limited to zero magnetic field due to the negative sign problem, a Majorana-based quantum Monte Carlo simulation has been performed for thermal transport in an effective model derived by a perturbation expansion with respect to the magnetic field, which avoids the sign problem by remaining within the Majorana fermion framework [54]. This study revealed a non-monotonic temperature dependence in $\kappa_{xy}/T$, featuring a weak peak at high temperature, but the result cannot explain the overall temperature dependence observed in experiments, including the overshooting behavior at low temperatures. For the regime beyond the perturbation, a continuous-time quantum Monte Carlo simulation in the original spin basis has been developed [55], but it also suffers from the negative sign problem at low temperatures and was not applied to the study of the thermal Hall transport.
@@
In this paper, we investigate the thermal Hall transport for an extended Kitaev model in a magnetic field, which includes the symmetric off-diagonal interactions, $\Gamma$ and $\Gamma^{\prime}$, using a finite-temperature tensor network method benchmarked with a thermal pure quantum state method. For comparison, we also perform classical Monte Carlo simulations in the classical limit of the model. We properly define the energy current from the model Hamiltonian and calculate the edge current in the systems with open boundaries. This treatment does not assume a specific origin for the thermal Hall transport, and includes all contributions in this spin model on an equal footing. We find that $\kappa_{xy}/T$ calculated for the pure Kitaev model shows a clear overshooting behavior, forming a hump at an intermediate temperature, which cannot be obtained by perturbation theory. We also investigate the effects of $\Gamma$ and $\Gamma^{\prime}$, showing that these interactions considerably affect the thermal Hall conductivity in both its magnitude and sign. In particular, we find that the thermal Hall conductivity is significantly enhanced by positive $\Gamma$ and negative $\Gamma^{\prime}$ in an intermediate temperature range. Furthermore, by comparing with the results of classical Monte Carlo simulations, we reveal the strong quantum nature of the $\Gamma$ contributions. Our comprehensive analyses suggest that these nonperturbative behaviors of the thermal Hall conductivity primarily stem from topological Majorana fermions. This insight would contribute to a deeper understanding of the experimental observations and offer a potential resolution to the existing discrepancies.
@@ II Model
Figure 1: (a) Schematic view of the extended Kitaev model on the honeycomb lattice in Eq. (2). Open circles represent magnetic ions hosting spin-$1/2$ degrees of freedom, while black dots denote nonmagnetic ligands forming an edge-sharing octahedral network. The $x$-, $y$-, and $z$-bonds in the Hamiltonian in Eq. (2) are indicated by green, blue, and red lines, respectively. The spin axes $(S^{x},S^{y},S^{z})$ are also shown, along with the crystallographic directions [$11\bar{2}$], [$\bar{1}10$], and [$111$]. (b) A honeycomb lattice cluster with $(L,L^{\prime})=(6,6)$, consisting of 72 sites, which is typically used in the XTRG calculations. We impose the periodic and open boundary conditions along the vertical and horizontal directions, respectively. We consider the thermal current $\bm{J}_{\mathrm{E}}$ in a downward direction indicated by the purple arrow. Dotted boxes represent the zigzag chains on which the energy current $J_{{\rm E},l}^{\parallel}$ in Eq. (9) is calculated. In the tensor network simulation, we consider a snake-like matrix product operator indicated by the blue-gray thick line behind the lattice.
@@
In the following sections, we consider the ferromagnetic Kitaev interaction by setting $K=-1$, which naturally arises from the exchange process in the $t_{2g}^{5}$ systems [14]. For $\Gamma$ and $\Gamma^{\prime}$, we consider both positive and negative cases to clarify their effects comprehensively. We take $k_{\rm B}=\hbar=1$ and set the length of the primitive translation vector of the honeycomb lattice to the unit length. We typically compute physical quantities for the model on a finite-size cluster with $(L,L^{\prime})=(6,6)$ shown in Fig. 1(b), including $N_{\rm s}=2\times L\times L^{\prime}=72$ spins.
@@ III.1 Thermal Hall conductivity
where $\langle\cdots\rangle_{T}$ represents the thermal average at temperature $T$. The coefficient $2/L^{\prime}$ is required as the thermal Hall conductivity is defined for the energy current per unit area. Note that Eq. (10) assumes $J_{\mathrm{E}}^{\parallel}$ is well localized at the edges.
@@ III.2 Tensor network method
As the initial condition of the XTRG method, we prepare $\rho(\beta_{0})$ with $\beta_{0}=10^{-7}$ through the approximated form $\rho(\beta_{0})\simeq 1-\beta_{0}\mathcal{H}$, where the Hamiltonian is represented as an MPO. We calculate the expectation value of an operator ${O}$ through thus obtained $\rho(\beta)$ as $\langle{O}\rangle=\mathrm{Tr}[{O}\rho]/\mathrm{Tr}\rho$ (note that the density matrix is not normalized). The temperature derivative of $\langle{O}\rangle$ is computed by numerical differentiation. In the following, we mainly show the data with $D=500$ for the $(L,L^{\prime})=(6,6)$ cluster. Benchmark calculations for the values of $D$ and the cluster size and geometry are shown in Appendix A.
@@ III.4 Classical Monte Carlo simulation
Finally, we describe the method of classical Monte Carlo simulation. In this method, an $S=1/2$ spin at each site is regarded as a classical vector with length $1/2$. The classical spin at site $i$ is parameterized by $\theta_{i}$ and $\phi_{i}$ as $\bm{S}_{i}=\frac{1}{2}(\sin\theta_{i}\cos\phi_{i},\sin\theta_{i}\sin\phi_{i},\cos\theta_{i})$. In the calculations of the thermal average, the integral $\int\prod_{i}d\phi_{i}d\theta_{i}\sin\theta_{i}$ is evaluated using the Markov-chain Monte Carlo method. To accelerate computations and avoid trapping spin configurations in local minima, we use the replica exchange method [91]. In the simulations, we prepare 48 replicas with different temperatures. We perform 10 000 000 MC steps for measurements after 10 000 MC steps for thermalization in the 800-site cluster with $(L,L^{\prime})=(10,40)$. The temperature derivative of $\langle O\rangle$ is evaluated by the correlation with ${\cal H}$ as
@@ IV.1.1 Magnetic field along [111] direction
We begin with the model with only the ferromagnetic Kitaev interaction ($K=-1$ and $\Gamma=\Gamma^{\prime}=0$) in an applied magnetic field along the $[111]$ direction ($h^{x}=h^{y}=h^{z}=h/\sqrt{3}$). In this setup, we expect $\kappa_{xy}/T$ to be quantized at a positive value of $\pi/12$ for small magnetic fields in the zero temperature limit [9]. Figures 5(a)-5(c) show the temperature dependence of the specific heat $C$, the magnetization along the magnetic field $M_{\parallel}$, and the flux density $W$, respectively given by
@@
Figure 6 also plots the contributions from each zigzag chain, $\langle J_{\mathrm{E},l}^{\parallel}\rangle_{T}$. We observe that these contributions exhibit similar temperature dependence, while they decrease rapidly with increasing distance from the edge, $l$. To highlight this behavior more clearly, we display the $l$ dependence of the energy current amplitude at several magnetic fields for $T\simeq 0.0186$ in Fig. 7. We find that the amplitude of the energy current is largest at the edge ($l=1$), and decreases almost exponentially toward the center of the system ($l=6$) [see Fig. 1(b)]. This result demonstrates that the energy current is dominated by the edge current, similar to the chiral edge current originating from the topologically-nontrivial Majorana band structure predicted by the perturbation theory. The contributions become negligibly small near the center, suggesting that the employed system size, $L=6$, is large enough to capture the essential features of this edge current. This also supports the validity of the assumption made in Eq. (10).
@@
Finally, let us discuss the asymptotic half-integer quantization at low temperatures. We note that while $\kappa_{xy}/T$ takes a value around the half-integer quantization $\pi/12$ at the lowest temperature calculated, it does not exhibit clear convergence to this value, even when the magnetic field is sufficiently small. This is presumably due to multiple reasons, including the finite bond dimension $D$, finite-size effects, and insufficiently low temperatures. First, as discussed for the specific heat above, the bond dimension $D=500$ is not sufficient to yield reliable results at low temperatures for $h\lesssim 0.04$. Second, finite-size effects become more pronounced at smaller $h$, where the excitation gap is expected to be smaller. This leads to a broader spatial extension of the edge current into the bulk  causing overlap between contributions from both edges in the finite-size cluster and thereby hampering accurate estimation of the edge current. Moreover, the finite length of the system along the edge direction may induce a gap in the chiral edge mode, further contributing to deviations from the quantization value. Lastly, asymptotic quantization is expected only at temperatures much lower than the excitation gap of topological quasiparticles. However, the precise value of this gap is not known beyond perturbation theory. Addressing these issues requires more sophisticated numerical methods capable of accessing larger-size systems at lower temperatures.
@@ IV.1.2 Field-direction dependence
Figure 8: Temperature dependence of (a) the specific heat, (b) the magnetization, (c) the flux density, (d) $\kappa_{xy}/T$, and (e) three-spin and (f) two-spin contributions to $\kappa_{xy}/T$ of the pure Kitaev model under magnetic fields parallel to $[111]$, $[11\bar{2}]$, and $[1\bar{1}0]$ with $h=0.04$ and $h=0.08$.
@@
Figure 8 shows the results for $h=0.04$ and $h=0.08$. We find that the specific heat and the magnetization exhibit no significant changes across different field directions, as shown in Figs. 8(a) and 8(b), respectively. In contrast, the flux density shows a significant decrease for the [$\bar{1}$10] direction, as shown in Fig. 8(c). This behavior may be associated with the presence of low-energy excitations, consistent with the prediction from the perturbation theory that the Majorana excitation gap does not open for a magnetic field applied in this direction [9].
The thermal Hall conductivity exhibits strong dependence on the direction of the applied magnetic field, consistent with predictions from the perturbation theory. As shown in Fig. 8(d), we find that $\kappa_{xy}/T$ reverses its sign from positive to negative when the field direction is changed from [$111$] to [$11\bar{2}$]. We also find that the three-spin contribution $\kappa_{xy}^{(L)}$ dominates in both cases and is responsible for the observed sign change of the total $\kappa_{xy}$. In addition, $\kappa_{xy}/T$ is vanishingly small for the [$\bar{1}11$] field. These results suggest that the topological Majorana fermion picture provides a plausible explanation for the field-direction dependence of the thermal Hall effect, even in the wide range of magnetic field strengths beyond the perturbative region.
We again emphasize that these behaviors persist even at field strengths beyond the critical field value, $h_{c}\simeq 0.024$. Although results based on the conventional magnon picture are not readily available for the [$11\bar{2}$] and [$\bar{1}11$] field directions, our findings consistently support the relevance of topological Majorana fermions in the enhanced thermal Hall response, even within the polarized regime beyond the quantum critical point.
We note that the vanishing of $\kappa_{xy}/T$ for the [$\bar{1}10$] field at all temperatures can be understood from the symmetry of the Hamiltonian. The Hamiltonian in Eq. (2) is invariant under the following combined operations: a $C_{2}$ rotation in spin space about the $[\bar{1}10]$ axis, which transforms $(S^{x},S^{y},S^{z})$ to $(-S^{y},-S^{x},-S^{z})$, and a rotation of the honeycomb plane in real space about the $z$ bond, which exchanges the $x$ and $y$ bonds. This real-space rotation of the honeycomb plane inverts the component of the position vector $\bm{r}$ along the zigzag edge, and thereby, the thermal current along the zigzag edge should be zero. Note that this symmetry argument remains valid even in the presence of the $\Gamma$ and $\Gamma^{\prime}$ interactions, as well as in the classical case.
@@ IV.2.1 Symmetric off-diagonal interaction $\Gamma^{\prime}$
First, we consider the effects of the $\Gamma^{\prime}$ interaction. Figure 9 displays the representative results at $h=0.04$ for $-0.02\leq\Gamma^{\prime}\leq 0.02$. Here, we set $\Gamma=0$ to isolate the influence of the $\Gamma^{\prime}$ term. As shown in Fig. 9(a), $\Gamma^{\prime}$ does not significantly change the overall temperature dependence of the specific heat, although the low-temperature peak shifts slightly to higher (lower) temperatures for negative (positive) $\Gamma^{\prime}$. We also observe that the magnetization increases (decreases) monotonically with negative (positive) $\Gamma^{\prime}$, as shown in Fig. 9(b). Compared with the results in Figs. 5(a) and 5(b), this behavior can be interpreted by considering $\Gamma^{\prime}$ as an effective magnetic field, whose strength increases (decreases) for a negative (positive) $\Gamma^{\prime}$. This interpretation appears to be consistent with the perturbation theory, which predicts that the Majorana band gap is proportional to $-\Gamma^{\prime}/|K|$ [56]. In contrast, as shown in Fig. 9(c), the flux density does not show considerable changes as a function of $\Gamma^{\prime}$, suggesting limitations of the effective field picture.

### v2

@@ I Introduction
where $k_{\mathrm{B}}$ and $\hbar$ denote the Boltzmann and Dirac constants, respectively. Here, $\nu$ takes $\pm 1$ depending on the direction of the magnetic field. The coefficient of $1/2$ is the consequence of the fractionalization; each Majorana fermion carries half the degrees of a conventional Dirac fermion. Thus, this quantized value is half of that for electrons with the same band topology, and hence, it is called the half-integer quantization. Detecting this peculiar behavior offers direct evidence of the Majorana fermions in the Kitaev QSL.
@@
For the Kitaev model without other additional interactions, efficient numerical simulation methods have been developed for studying thermodynamic properties and spin dynamics at finite temperatures, utilizing the Majorana fermion representation [49, 50, 51, 52, 53]. This approach explicitly leverages the emergent quasiparticle picture in fractionalization. Although these methods are limited to zero magnetic field due to the negative sign problem, a Majorana-based quantum Monte Carlo simulation has been performed for thermal transport in an effective model derived by a perturbation expansion with respect to the magnetic field, which avoids the sign problem by remaining within the Majorana fermion framework [54]. This study revealed a non-monotonic temperature dependence in $\kappa_{xy}/T$, featuring a weak peak at high temperature, but the result cannot explain the overall temperature dependence observed in experiments, including the overshooting behavior at low temperatures. For the regime beyond the perturbation, a continuous-time quantum Monte Carlo simulation in the original spin basis has been developed [55], but it also suffers from the negative sign problem at low temperatures and was not applied to the study of the thermal Hall transport. In addition, because the standard Kubo formula based on the linear-response theory requires information on dynamical physical quantities, reliable numerical calculations of thermal transport for microscopic models are challenging.
@@
In this paper, we investigate the thermal Hall transport for an extended Kitaev model in a magnetic field, which includes the symmetric off-diagonal interactions, $\Gamma$ and $\Gamma^{\prime}$, using a finite-temperature tensor network method benchmarked with a thermal pure quantum state method. For comparison, we also perform classical Monte Carlo simulations in the classical limit of the model. We properly define the energy current from the model Hamiltonian and calculate the edge current in the systems with open boundaries. This treatment does not assume a specific origin for the thermal Hall transport, and includes all contributions in this spin model on an equal footing. Furthermore, this approach allows us to evaluate the thermal Hall conductivity without the explicit computation of dynamical quantities. We find that $\kappa_{xy}/T$ calculated for the pure Kitaev model shows a clear overshooting behavior, forming a hump at an intermediate temperature, which cannot be obtained by perturbation theory. We also investigate the effects of $\Gamma$ and $\Gamma^{\prime}$, showing that these interactions considerably affect the thermal Hall conductivity in both its magnitude and sign. In particular, we find that the thermal Hall conductivity is significantly enhanced by positive $\Gamma$ and negative $\Gamma^{\prime}$ in an intermediate temperature range. Furthermore, by comparing with the results of classical Monte Carlo simulations, we reveal the strong quantum nature of the $\Gamma$ contributions. Our comprehensive analyses suggest that these nonperturbative behaviors of the thermal Hall conductivity primarily stem from topological Majorana fermions. This insight would contribute to a deeper understanding of the experimental observations and offer a potential resolution to the existing discrepancies.
@@ II Model
[…]
@@
In the following sections, we consider the ferromagnetic Kitaev interaction by setting K=1K=-1, which naturally arises from the exchange process in the $t_{2g}^{5}$ systems [14]. For $\Gamma$ and $\Gamma^{\prime}$, we consider both positive and negative cases to clarify their effects comprehensively. We take $k_{\rm B}=\hbar=1$ and set the length of the primitive translation vector of the honeycomb lattice to the unit length. We typically compute physical quantities for the model on a finite-size cluster with $(L,L^{\prime})=(6,6)$ shown in Fig. 1(b), including $N_{\rm s}=2\times L\times L^{\prime}=72$ spins.
@@ III.1 Thermal Hall conductivity
where $\langle\cdots\rangle_{T}$ represents the thermal average at temperature $T$. The coefficient $2/L^{\prime}$ is required as the thermal Hall conductivity is defined for the energy current per unit area. Note that Eq. (10) assumes $J_{\mathrm{E}}^{\parallel}$ is well localized at the edges. This definition of the thermal Hall conductivity is similar to the one used in Ref. [80]. As noted in the literature, the half-quantization of $\kappa_{xy}/T$ is expected to occur within the temperature range of $v/L^{\prime}\ll T\ll\Delta_{\mathrm{bulk}}$, where $v$ is the velocity of the edge mode and $\Delta_{\mathrm{bulk}}$ is the excitation gap of topological quasiparticles. We will return to this point in Sec. IV.1.1.
@@ III.2 Tensor network method
As the initial condition of the XTRG method, we prepare $\rho(\beta_{0})$ with $\beta_{0}=10^{-7}$ through the approximated form $\rho(\beta_{0})\simeq 1-\beta_{0}\mathcal{H}$, where the Hamiltonian is represented as an MPO. We calculate the expectation value of an operator O{O} through thus obtained $\rho(\beta)$ as $\langle{O}\rangle=\mathrm{Tr}[{O}\rho]/\mathrm{Tr}\rho$ (note that the density matrix is not normalized). The temperature derivative of $\langle{O}\rangle$ is computed by numerical differentiation. In the following, we mainly show the data with $D=500$ for the $(L,L^{\prime})=(6,6)$ cluster. Benchmark calculations for the values of $D$ and the cluster size and geometry are shown in Appendix A.
@@ III.4 Classical Monte Carlo simulation
[…]
@@ IV.1.1 Magnetic field along [111] direction
We begin with the model with only the ferromagnetic Kitaev interaction (K=1K=-1 and $\Gamma=\Gamma^{\prime}=0$) in an applied magnetic field along the $[111]$ direction ($h^{x}=h^{y}=h^{z}=h/\sqrt{3}$). In this setup, we expect $\kappa_{xy}/T$ to be quantized at a positive value of $\pi/12$ for small magnetic fields in the zero temperature limit [9]. Figures 5(a)-5(c) show the temperature dependence of the specific heat $C$, the magnetization along the magnetic field $M_{\parallel}$, and the flux density $W$, respectively given by
@@
Figure 6 also plots the contributions from each zigzag chain, $\langle J_{\mathrm{E},l}^{\parallel}\rangle_{T}$. We observe that these contributions exhibit similar temperature dependence, while they decrease rapidly with increasing distance from the edge, $l$. To highlight this behavior more clearly, we display the $l$ dependence of the energy current amplitude at several magnetic fields for $T\simeq 0.0186$ in Fig. 7. We find that the amplitude of the energy current is largest at the edge ($l=1$), and decreases almost exponentially toward the center of the system ($l=6$) [see Fig. 1(b)]. This result demonstrates that the energy current is dominated by the edge current, similar to the chiral edge current originating from the topologically-nontrivial Majorana band structure predicted by the perturbation theory. The contributions become negligibly small near the center, suggesting that the employed system size, $L=6$, is large enough to capture the essential features of this edge current. Surprisingly, similar exponential decays are also observed at higher temperatures as shown in Appendix D. This supports the assumption made in Eq. (10), across a broad range of temperatures and magnetic fields in our calculations.
@@
Finally, let us discuss the asymptotic half-integer quantization at low temperatures. We note that while $\kappa_{xy}/T$ takes a value around the half-integer quantization $\pi/12$ at the lowest temperature calculated, it does not exhibit clear convergence to this value, even when the magnetic field is sufficiently small. This is presumably due to multiple reasons, including the finite bond dimension $D$, finite-size effects, and insufficiently low temperatures. First, as discussed for the specific heat above, the bond dimension $D=500$ is not sufficient to yield reliable results at low temperatures for $h\lesssim 0.04$. Second, finite-size effects become more pronounced at smaller $h$, where the excitation gap is expected to be smaller. This leads to a broader spatial extension of the edge current into the bulk, causing overlap between contributions from both edges in the finite-size cluster and thereby hampering accurate estimation of the edge current. Moreover, the finite length of the cluster along the edges, $L^{\prime}$, further contributes to deviations from quantization by discretizing the chiral edge mode; clear quantization is expected only when $T\gg v/L^{\prime}$ [80]. Lastly, asymptotic quantization is expected only at temperatures much lower than the topological bulk gap, i.e., when $T\ll\Delta_{\mathrm{bulk}}$ [80]. However, the precise value of this gap is not known beyond perturbation theory. Addressing these issues requires more sophisticated numerical methods capable of accessing larger-size systems at lower temperatures.
@@ IV.1.2 Field-direction dependence
Figure 8: Temperature dependence of (a) the specific heat, (b) the magnetization, (c) the flux density, (d) $\kappa_{xy}/T$, and (e) three-spin and (f) two-spin contributions to $\kappa_{xy}/T$ of the pure Kitaev model under magnetic fields parallel to $[111]$, $11\bar{2}]$, and $1\bar{1}0]$ with $h=0.04$ and $h=0.08$.
@@
Figure 8 shows the results for $h=0.04$ and $h=0.08$. We find that the specific heat and the magnetization exhibit no significant changes across different field directions, as shown in Figs. 8(a) and 8(b), respectively. In contrast, the flux density shows a significant decrease for the $\bar{1}$10] direction, as shown in Fig. 8(c). This behavior may be associated with the presence of low-energy excitations, consistent with the prediction from the perturbation theory that the Majorana excitation gap does not open for a magnetic field applied in this direction [9].
The thermal Hall conductivity exhibits strong dependence on the direction of the applied magnetic field, consistent with predictions from the perturbation theory. As shown in Fig. 8(d), we find that $\kappa_{xy}/T$ reverses its sign from positive to negative when the field direction is changed from [$111$] to $11\bar{2}$]. We also find that the three-spin contribution $\kappa_{xy}^{(L)}$ dominates in both cases and is responsible for the observed sign change of the total $\kappa_{xy}$. In addition, $\kappa_{xy}/T$ is vanishingly small for the $\bar{1}11$] field. These results suggest that the topological Majorana fermion picture provides a plausible explanation for the field-direction dependence of the thermal Hall effect, even in the wide range of magnetic field strengths beyond the perturbative region.
We again emphasize that these behaviors persist even at field strengths beyond the critical field value, $h_{c}\simeq 0.024$. Although results based on the conventional magnon picture are not readily available for the $11\bar{2}$] and $\bar{1}11$] field directions, our findings consistently support the relevance of topological Majorana fermions in the enhanced thermal Hall response, even within the polarized regime beyond the quantum critical point.
We note that the vanishing of $\kappa_{xy}/T$ for the $\bar{1}10$] field at all temperatures can be understood from the symmetry of the Hamiltonian. The Hamiltonian in Eq. (2) is invariant under the following combined operations: a $C_{2}$ rotation in spin space about the $\bar{1}10]$ axis, which transforms $(S^{x},S^{y},S^{z})$ to $(-S^{y},-S^{x},-S^{z})$, and a rotation of the honeycomb plane in real space about the $z$ bond, which exchanges the $x$ and $y$ bonds. This real-space rotation of the honeycomb plane inverts the component of the position vector $\bm{r}$ along the zigzag edge, and thereby, the thermal current along the zigzag edge should be zero. Note that this symmetry argument remains valid even in the presence of the $\Gamma$ and $\Gamma^{\prime}$ interactions, as well as in the classical case.
@@ IV.2.1 Symmetric off-diagonal interaction $\Gamma^{\prime}$
First, we consider the effects of the $\Gamma^{\prime}$ interaction. Figure 9 displays the representative results at $h=0.04$ for $0.02\leq\Gamma^{\prime}\leq 0.02$. Here, we set $\Gamma=0$ to isolate the influence of the $\Gamma^{\prime}$ term. As shown in Fig. 9(a), $\Gamma^{\prime}$ does not significantly change the overall temperature dependence of the specific heat, although the low-temperature peak shifts slightly to higher (lower) temperatures for negative (positive) $\Gamma^{\prime}$. We also observe that the magnetization increases (decreases) monotonically with negative (positive) $\Gamma^{\prime}$, as shown in Fig. 9(b). Compared with the results in Figs. 5(a) and 5(b), this behavior can be interpreted by considering $\Gamma^{\prime}$ as an effective magnetic field, whose strength increases (decreases) for a negative (positive) $\Gamma^{\prime}$. This interpretation appears to be consistent with the perturbation theory, which predicts that the Majorana band gap is proportional to $\Gamma^{\prime}/|K|$ [56]. In contrast, as shown in Fig. 9(c), the flux density does not show considerable changes as a function of $\Gamma^{\prime}$, suggesting limitations of the effective field picture.